Calculate the number of neutrons in 100g of Mohr salt.

Grade:11

1 Answers

divyanshu
6 years ago
Mohr’s salt is (NH4)2Fe(SO42· 6H2O
It has N=2 atoms, H=20 atoms, Fe=1 atom, S= 2 atoms and O= 14 atoms in a single molecule.
Therefore, it’s mass is 14×2 + 20 + 56 + 32×2 + 16×14= 392 grams/mol.
Thus 100 grams is 100/392 mols of Mohr’s salt. which is equal to 0.255 mols.
Which has 0.255× NA(6.022× 1023) number of molecules.
Each molecule has 7*2 + 0 + 30 + 16*2 + 8*14 neutrons, which is equal to 188 neutrons.
 
Thus, the total number of neutrons in 100 grams of Mohr’s salt is 188*0.255* NA = 2.887 × 1025.
 
The Uber Nerd.